OpenStudy (anonymous):

determine whether the sequence an=2e^(2n/(n+2)) converges or diverges. if it converges, find the limit

OpenStudy (anonymous):

need help

jimthompson5910 (jim_thompson5910):

Focus on 2n/(n+2) for a moment

jimthompson5910 (jim_thompson5910):

what happens as n goes off to infinity?

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

it increases?

jimthompson5910 (jim_thompson5910):

if n = 1000, then what is 2n/(n+2) equal to?

OpenStudy (anonymous):

it goes to infinty and beyond

OpenStudy (anonymous):

its 1.996

OpenStudy (anonymous):

a very small number...

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ok now what is 2n/(n+2) equal to when n = 10,000 ?

OpenStudy (anonymous):

an even smaller number

OpenStudy (freckles):

\[\frac{2n}{n+2} \approx \frac{2n}{n} \text{ for \large } n \]

OpenStudy (anonymous):

1.999600000 to be exact

OpenStudy (freckles):

do you know where 2n/n goes as n continues to get bigger?

OpenStudy (anonymous):

i dont know

OpenStudy (freckles):

ok just do it @jim_thompson5910 's way then

jimthompson5910 (jim_thompson5910):

keep going, what about if n = 100,000 ?

OpenStudy (anonymous):

do you want me to do the 2n/n or 2n/n+2?

jimthompson5910 (jim_thompson5910):

let's use the original

OpenStudy (anonymous):

i got 1.999960001

jimthompson5910 (jim_thompson5910):

and if n = 1,000,000 then 2n/(n+2) = 1.999996000 and if n = 10,000,000 then 2n/(n+2) = 1.999999600 and if n = 100,000,000 then 2n/(n+2) = 1.999999960 and so on... so if you keep bumping up n larger and larger, you'll find that 2n/(n+2) approaches 2 the leading terms 2n and n divide to get 2n/n = 2

OpenStudy (anonymous):

oh alright

jimthompson5910 (jim_thompson5910):

so back to 2e^(2n/(n+2))

jimthompson5910 (jim_thompson5910):

what happens to 2e^(2n/(n+2)) as n --> infinity?

OpenStudy (anonymous):

it approaches the number 15

jimthompson5910 (jim_thompson5910):

what does 2e^(2n/(n+2)) simplify to when n approaches infinity? give the exact form of it

OpenStudy (anonymous):

u mean like use the L'hospital rule

jimthompson5910 (jim_thompson5910):

let's say I had sin(x+2) and I wanted to know the exact value of that expression above when x = 4. I just plug in x = 4 and try to simplify as much as possible sin(x+2) = sin(4+2) = sin(6) sin(6) can't be simplified further so I leave it as it is

jimthompson5910 (jim_thompson5910):

sin(6) has an approximate decimal form, but sometimes it's best to leave it in exact form

OpenStudy (anonymous):

in this case we have 2e^2n/n+2 which simplifies to 2lne^2n/n+2 = 4n/n+2

jimthompson5910 (jim_thompson5910):

no, we already know 2n/(n+2) approaches 2 as n approaches infinity so just replace 2n/(n+2) with 2 and you're done \[\LARGE \lim_{n \to \infty}\left(2e^{\frac{2n}{n+2}}\right) = 2e^{2}\]

jimthompson5910 (jim_thompson5910):

so the whole thing approaches 2e^2 as n approaches infinity

OpenStudy (anonymous):

so the sequence converges

jimthompson5910 (jim_thompson5910):

yes because 2n/(n+2) converged

OpenStudy (anonymous):

@freckles you still here?

OpenStudy (anonymous):

@jim_thompson5910 i got a quick question

OpenStudy (anonymous):

so you would do every exponent problem like this for example if we had 3 instead of 2 it would be 2e^3

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

what is the full problem

OpenStudy (anonymous):

no just asking if we had another exponent problem like this one

jimthompson5910 (jim_thompson5910):

yes you break it up like that

OpenStudy (anonymous):

okay and what about when its a natural log problem you break it up too?

jimthompson5910 (jim_thompson5910):

yes if you had ln( 2n/(n+2) ) you would focus on the 2n/(n+2) first and see that it approaches 2 so ln( 2n/(n+2) ) approaches ln( 2 )

OpenStudy (anonymous):

okay thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome